OpenStudy (anonymous):

2(7-x)divided by 5=x how do i simplfy and solve this problem

OpenStudy (kewlgeek555):

Okay, is this the equation?: \[2(7-x) \div 5=x\]

OpenStudy (anonymous):

\[2(7-x)5=x\]

OpenStudy (kewlgeek555):

Okay, so its 2 times (7-x) times 5 = x?

OpenStudy (kewlgeek555):

First, let's work on the left hand side of your equation, the 2 * (7-x) * 5 This means, for instance, to see if it can be simplified at all. 7-x evaluates to 7-x Multiply 2 by 7-x we multiply 2 by each term in 7-x term by term. This is the distributive property of multiplication. Multiply 2 and 7 1 2 × 7 = 14 Multiply 2 and -x

OpenStudy (kewlgeek555):

That will be 2x

OpenStudy (kewlgeek555):

Now your equation is 14 - 2x * 5

OpenStudy (kewlgeek555):

Multiply 14-2x by 5 we multiply 14-2x by each term in 5 term by term. This is the distributive property of multiplication. Multiply 14 and 5 1 14 × 5 = 70

OpenStudy (kewlgeek555):

So now your equation is 70-10x